# ProGeny I: a new simple stellar population generator and impact of isochrones / stellar atmospheres / initial mass functions

### v2

@@ 3.2.1 Base and Extended Stellar Templates
Within ProGeny we include a dense grid of ‘base’ templates that cover the main sequence of most stars (see Table 3). These were noted to not extend high enough in temperature so smoothly connect with the ‘hot’ stellar atmospheres used (discussed below). To improve coverage ProGeny has the capability to supplement any ‘base’ templates with an extension from a different library (‘extend’). When the SSP is being generated the ‘base’ templates are prioritised when matching parameters, with a user-controllable parameter specifying how much closer matched in temperature, logG and metallicity space the ‘extend’ atmosphere has to be before it is used instead.
@@
The result is that the ‘base’ and ‘extend’ templates generally cover the vast majority of the lifetime of stars for any typical IMF regardless of the isochrones used. This is clear when considering Figure 3 where we display all stellar atmosphere library grid points.
@@
Figure 3: Coverage of the fiducial stellar atmospheres used in ProGeny with random samples from the MIST / PARSEC and BaSTI isochrones overlaid. The density of isochrone points is broadly indicative of which stellar atmospheres are the most important, and where most star light is coming from over the lifetime of stellar evolution. The small vertical dithering for ‘base’ and ‘extend’ templates are to reflect the presence of metallicity (Z) gridding (this is not available for ‘hot’ or ‘AGB’ stellar atmospheres).
@@
For good all-round coverage in the dimensions of temperature, effective surface gravity (logG) and metallicity (Z) we use the C3K atmospheric spectra included with FSPS as our fiducial ‘base’ spectra (Conroy et al., 2018) . This covers temperatures from 2,000 to 50,000 K; logG from -1 to 5.5; and metallicity from -2.1 to0.5. Compared to alternatives, C3K has very broad coverage in temperature and logG, but is a bit more limited in terms of the available metallicities.
@@
The fiducial ‘extend’ template spectra comes from the Allard et al. (2012) version of PHOENIX spectra777available at archive.stsci.edu/hlsps/reference-atlases/cdbs/grid/phoenix/. For details on the PHOENIX stellar atmosphere generation software see Hauschildt et al. (1999) and associated works. Currently this library covers 2,000 – 70,000 K, which is reasonably hot, but means we do still need a dedicated ‘hot’ atmosphere library too to achieve the appropriate coverage of remnant stars when using the MIST isochrones (especially white dwarfs). If using the PARSEC isochrones then the combination of ‘base’ and ‘extend’ libraries achieves good coverage of all but the AGB evolutionary phases. The Allard library covers logG from 0 to 5.5; and metallicity from -4 to 0.5, giving it better metallicity coverage than C3K, but not extending as far into the low logG AGB stellar atmospheres. Using C3K (as ‘base’) and Allard (as ‘extend’) gives excellent all-round parameter coverage for most phases of stellar evolution, but with additional treatment for AGB phases and ultra hot stars required (see below).
@@
An alternative ‘base’ template spectra comes from the Husser et al. (2013) Gottingen version of the PHOENIX spectra888available at ftp://phoenix.astro.physik.uni-goettingen.de/HiResFITS/. We used the log10($\alpha$) = 0 (not enhanced) version of the library. The high resolution spectra (with better wavelength coverage than the mid resolution also available) was then re-binned to match the sampling resolution of other ProGeny stellar atmospheres (which is the same as the CB19 high resolution library by default). Currently this library only extends to 12, [^0] K, where the website claims the plan is to eventually extend it to 25, [^0] K. Given the passage of time since the library’s release (11 years) it appears this extension may never happen. This necessitates the use of a broader range ‘extend’ library when using the Husser stellar atmospheres within ProGeny, and is the reason we do not use the Husser library as our fiducial ‘base’ template spectra.
@@
Figure 4: Colour-colour comparison of different stellar atmospheres as a function of age. In all cases the ‘extend’ library used is Allard PHOENIX (i.e. the blue PHOENIX label corresponds to the ‘base’ Husser PHOENIX variant). Other versions of this Figure (for different metallicities) can be viewed at rpubs.com/asgr.
@@
Figure 4 presents the colour – colour (NUV - r versus r - Ks) scatter between different combinations of stellar atmospheres of table 3 when using the same isochrone (MIST), IMF (Chabrier) and metallicity (solar). The variations peak for r - Ks colour at 10 Myr ($\sim 0.5$ mag scatter). Beyond 100 Myr the colour scatter is $\sim 0.1$ mag. Since only very young stellar populations have notable colour scatter due to the choice of atmosphere, we can generally assume the stellar mass is only marginally affected by this decision. Even the star formation rate (the integral of stellar mass formed up to 100 Myr in ProSpect) will see little impact since it is the NUV - r axis (with very little scatter) the provides most of the constraining information when fitting a galaxy SED.
@@ 3.2.2 TP-AGB and AGB Stellar Templates
A key extension to the ‘base’ and the ‘extend’ stellar atmospheres is the complex and highly variable asymptotic giant branch (AGB) phases of stellar evolution. These are both highly complex to model theoretically, and are generally not well understood in terms of the stellar atmosphere spectra produced. In temporal order these phases cover the pre-AGB, early-AGB, AGB (including TP-AGB) and post-AGB periods of stellar evolution. The reason for the modelling difficulty is due to the rapid timescales of variability, and the complex interplay between the different layers of the stars (which both influences the behaviour of evolutionary tracks and the predicted spectra of the stellar atmospheres). A number of works have focussed on the complexity of this phase of stellar evolution, and there is a general consensus that for certain periods of stellar evolution (a few Gyr in age) the contribution can be highly significant due to the sheer size and luminosity of such stars, even though the amount of stellar mass involved might be small (see Maraston, 2005).
@@
As noted above, different isochrone models have quite different predictions for the AGB phases (see Figure 1). Common to other works, we choose an empirical library of AGB spectra since there is no theoretical template set that offers the AGB parameter coverage required for constructing an SSP. Specifically, we use the small template library released in Lançon & Mouhcine (2002). This contains time-averaged AGB spectra combined over a number of years of observations of the same AGB stars covering a reasonable range of AGB classes and temperatures (2,000 – 4,000 K). This template set is generally considered to be a good compromise for AGB spectra since it captures the temporally extreme pulsations that in practice vary even more rapidly than stellar evolution isochrones can describe. The Lancon AGB spectra was extended beyond the original 5,100 – 24,900 Å range to cover the total range needed for ProGeny using a blackbody extension (at the appropriate temperature) at either end.
@@
In terms of parameter coverage, the AGB stellar atmospheres have the sparsest and least regular coverage, with only 14 templates covering the approximate temperature range 2,000 – 4,000 K. They also do not have a specified (i.e. known) effective surface gravity (logG), so for the purposes of isochrone parameter interpolation they are placed at a fiducial value of -1 (as can be seen in Figure 3). The consequence of this is mitigated by the ‘base’ and ‘extend’ stellar atmospheres extending down to logG of 0, which covers the majority of AGB evolutionary phases in practice. It is only the most extreme TP-AGB pulsations that require the use of the empirical Lancon libraries provided here.
3.2.3 Hot and Remnant Stellar Templates
@@
The final stellar atmosphere variant that must be included for suitable isochrone coverage is ‘hot’ stars including stellar remnants (i.e. white dwarfs and neutron stars). This includes most stars hotter than 50,000 K and ones with an effective surface gravity (logG) 6 $\text{cm}/\text{s}^{2}$ or greater, i.e. extremely compact and hot objects. It is notable from 3 that only the MIST isochrones have significant coverage inside this regime, i.e. the PARSEC and BaSTI isochrones do not use or require ‘hot’ stellar atmospheres. The hot stellar atmospheres are important for producing at least some of the UV upturn for old stellar populations with ProGeny (see Figure 2), but it should be noted that there is still a great deal of uncertainty about what contributes to UV upturn features in elliptical galaxies. For instance, it is not well understood how dominant exposed core helium burning could be for producing UV photons, and the thorough modelling of binary evolution certainly has a significant effect (Stanway & Eldridge, 2018).
@@
The fiducial ‘hot’ template spectra comes from Werner et al. (2003) (and earlier works). It is a combination of the TMAP2 and TMAP3 Werner data available999svo2.cab.inta-csic.es/theory/newov2/index.php. The library used reaches up to 190, [^0] K, and covers the range of Teff and logG for the majority of hot stars and white dwarf remnants observed on the MIST isochrones (Teff 20,000 – 190,000 K and logG 5 – 9 $\text{cm}/\text{s}^{2}$). For this reason it works reasonably well for all hot phases (including helium burning) and remnants (including white dwarfs). The very coolest white dwarfs are not quite covered (below 20,000 K), but these barely exist within the current age of the Universe, and are such low luminosity that they contribute very little flux to any reasonable SSP.
@@ 3.3 Initial Mass Function
Figure 5: Comparison of the different classic IMFs available as standard within ProGeny. In all cases 1 solar mass of stars have been formed by the IMF. The Chabrier and Kroupa IMFs are in general very similar. Note the larger value of the low mass cut off seen for the Salpeter IMF — this is to stop too many extremely low mass stars being formed.
@@
In total, the pre-built ProGeny IMFs come in seven different flavours of IMF: Salpeter, Kroupa, Chabrier, age-evolving Kroupa, age-evolving Lacey, metallicity evolving Kroupa, and metallically evolving Lacey. By following the included examples it is relatively easy for users to encode their own Universal IMFs, or even other evolving (either age or metallicity) variants. An obvious extension might be to parameterise the evolution via a property other than stellar population age. It often makes more sense to think of IMF evolution as being relative to cosmic age rather than age of a particular stellar population (which can be achieved by specifying the look-back time of observation in ProGeny), or perhaps relative to the metallicity enrichment history of the star formation event in question. The basic age-evolving IMFs available with ProGeny are shown in Figure 6, but note the exact implementation presented here is somewhat arbitrary — there is no well informed model for exactly how the IMF might evolve with age (or metallicity), so the existence of these functions is really to aid user experimentation (i.e. the default values should not be treated as especially well justified). Longer term work (beyond the scope of this paper), now that ProGeny exists with this degree of flexibility, will be to better justify if (and how) these evolving IMF forms can be used for fitting SEDs with ProSpect.
Figure 6: Comparison of the different age-evolving IMFs available as standard within ProGeny. In all cases 1 solar mass of stars have been formed by the IMF.
@@ 4 Comparisons and Discussion
Below we make some instructive comparisons of ProGeny. First internally (where we vary the isochrones, stellar atmospheres and IMF), and secondly to well know literature SSPs (that are mixture of similar and dissimilar SPLs).
4.1 ProGeny Comparisons
@@
It is important to understand the role that the key components in ProGeny have when generating spectra at specific ages and metallicities (the core behaviour of any SSP). The most significant variations in spectral output come from the choice of isochrone (MIST, PARSEC, or BaSTI in the case of ProGeny) and the type of IMF (both fixed and evolving, in our case).
4.1.1 Internal Comparisons

### v3

@@ 3.2.1 Base and Extended Stellar Templates
Within ProGeny we include a dense grid of ‘base’ templates that cover the main sequence of most stars (see Table 3). These were noted to not extend high enough in temperature so smoothly connect with the ‘hot’ stellar spectra used (discussed below). To improve coverage ProGeny has the capability to supplement any ‘base’ templates with an extension from a different library (‘extend’). When the SSP is being generated the ‘base’ templates are prioritised when matching parameters, with a user-controllable parameter specifying how much closer matched in temperature, surface gravity and metallicity space the ‘extend’ stellar spectrum has to be before it is used instead.
@@
The result is that the ‘base’ and ‘extend’ templates generally cover the vast majority of the lifetime of stars for any typical IMF regardless of the isochrones used. This is clear when considering Figure 3 where we display all stellar spectrum library grid points.
@@
Figure 3: Coverage of the fiducial stellar spectra used in ProGeny with random samples from the MIST / PARSEC and BaSTI isochrones overlaid. The density of isochrone points is broadly indicative of which stellar spectra are the most important, and where most star light is coming from over the lifetime of stellar evolution. The small vertical dithering for ‘base’ and ‘extend’ templates are to reflect the presence of metallicity (Z) gridding (this is not available for ‘hot’ or ‘AGB’ stellar spectra).
@@
For good all-round coverage in the dimensions of temperature, surface gravity, and metallicity we use the C3K atmospheric spectra included with Flexible Stellar Population Synthesis as our fiducial ‘base’ spectra (FSPS; Conroy et al., 2018) . This covers temperatures from 2,000 to 50,000 K; surface gravity from -1 to 5.5; and metallicity from -2.1 to 0.5. Compared to alternatives, C3K has very broad coverage in temperature and surface gravity, but is a bit more limited in terms of the available metallicities. In general C3K offers good coverage for even our youngest OB stars, but we note that should isochrones be used in the future that have stellar evolution younger than 100,000 yr then additional spectral templates would be needed to cover more extreme massive/young OB star spectra.
@@
The fiducial ‘extend’ template spectra comes from the Allard et al. (2012) version of PHOENIX spectra888available at archive.stsci.edu/hlsps/reference-atlases/cdbs/grid/phoenix/. For general details on the PHOENIX stellar spectrum generation software see Hauschildt et al. (1999) and associated works. This library is often referred to as ‘BT-Settl’, but we refer to it as ‘Allard’ within ProGeny and in this paper. The focus of this library was for very low mass stars (including brown dwarfs), but it covers more typical stellar populations also. Currently the Allard library covers 2,000 – 70,000 K, which is reasonably hot (certainly hot enough for the youngest OB stars generated by any available ProGeny isochrones), but means we do still need a dedicated ‘hot’ stellar spectra library too to achieve the appropriate coverage of remnant stars (i.e. white dwarfs) when using the MIST isochrones. If using the MIST or PARSEC isochrones then the combination of ‘base’ and ‘extend’ libraries achieves good coverage of all but the AGB evolutionary phases. The Allard library covers surface gravity from 0 to 5.5; and metallicity from -4 to 0.5, giving it better metallicity coverage than C3K, but not extending as far into the low surface gravity AGB stellar spectra. Using C3K (as ‘base’) and Allard (as ‘extend’) gives excellent all-round parameter coverage for most phases of stellar evolution, but with additional treatment for AGB phases and ultra hot stars required (see below).
@@
An alternative ‘base’ template spectra comes from the Husser et al. (2013) Gottingen version of the PHOENIX spectra999available at ftp://phoenix.astro.physik.uni-goettingen.de/HiResFITS/. We used the log10($\alpha$/Fe) = 0 (not enhanced) version of the library. The high resolution spectra (with better wavelength coverage than the mid resolution also available) was then re-binned to match the sampling resolution of other ProGeny stellar spectra (which is the same as the CB19 high resolution library by default). Currently this library only extends to 12,000 K, where the website claims the plan is to eventually extend it to 25,000 K. Given the passage of time since the library’s release (11 years) it appears this extension may never happen. This necessitates the use of a broader range ‘extend’ library when using the Husser stellar spectra within ProGeny, and is the reason we do not use the Husser library as our fiducial ‘base’ template spectra.
One phase of stellar evolution that is not well captured with current default stellar templates in ProGeny is Wolf-Rayet (WR) stars. This extreme phase of stellar evolution occurs between 1–10 Myr in the most massive (tens of solar masses) stars. Temperatures can reach millions of Kelvin, and in MIST the hottest WR stars evolve up to 320,000 K. With the default ProGeny stellar spectral templates these hottest WR stars get represented with a blackbody spectrum, which is certainly not fully accurate. For better representing WR evolution additional spectral templates can be used (e.g. CMFGEN, PoWR; Hillier & Miller, 1998; Todt et al., 2015, respectively). Keen users can add their own spectral templates quite easily, but longterm we will aim to add appropriate WR templates.
@@
Figure 4: Colour-colour comparison of different stellar spectra as a function of age for Z using MIST isochrones. In all cases the ‘extend’ library used is Allard PHOENIX (i.e. the blue PHOENIX label corresponds to the ‘base’ Husser PHOENIX variant). Other versions of this Figure (for different metallicities and isochrones) can be viewed in Appendix B.
@@
Figure 4 presents the colour – colour (NUV - r versus r - Ks) scatter between different combinations of stellar spectra of table 3 when using the same isochrone (MIST), IMF (Chabrier) and metallicity (solar). The variations peak for r - Ks colour at 10 Myr ($\sim 0.5$ mag scatter). Beyond 100 Myr the colour scatter is $\sim 0.1$ mag. Since only very young stellar populations have notable colour scatter due to the choice of stellar spectrum, we can generally assume the stellar mass is only marginally affected by this decision. Even the star formation rate (the integral of stellar mass formed up to 100 Myr in ProSpect) will see little impact since it is the NUV - r axis (with very little scatter) the provides most of the constraining information when fitting a galaxy SED. We note that BC03 does show a bluer NUV - r colour for the oldest (10 Gyr) populations. This is reflective of it possessing a stronger UV-upturn for old stars, likely due it incorporating UV bright planetary nebulae spectra (this is noted as a private communication in Vazdekis et al., 2016).
@@ 3.2.2 TP-AGB and AGB Stellar Templates
A key extension to the ‘base’ and the ‘extend’ stellar spectra is the complex and highly variable asymptotic giant branch (AGB) phases of stellar evolution. These are both highly complex to model theoretically, and are generally not well understood in terms of the stellar spectra produced. In temporal order these phases cover the pre-AGB, early-AGB, AGB (including TP-AGB) and post-AGB periods of stellar evolution. The reason for the modelling difficulty is due to the rapid timescales of variability, and the complex interplay between the different layers of the stars (which both influences the behaviour of evolutionary tracks and the predicted spectra of the stellar spectra). A number of works have focussed on the complexity of this phase of stellar evolution, and there is a general consensus that for certain periods of stellar evolution (a few Gyr in age) the contribution can be highly significant due to the sheer size and luminosity of such stars, even though the amount of stellar mass involved might be small (see Maraston, 2005; Conroy et al., 2010, and discussions therein).
@@
As noted above, different isochrone models have quite different predictions for the AGB phases (see Figure 1). Common to other works, we choose an empirical library of AGB spectra since there is no theoretical template set that offers the AGB parameter coverage required for constructing an SSP. Specifically, we use the small template library released in Lançon & Mouhcine (2002) as first used for SSP generation in Maraston (2005) and later in Conroy et al. (2010). This contains time-averaged AGB spectra combined over a number of years of observations of the same AGB stars covering a reasonable range of AGB classes and temperatures (2,000 – 4,000 K). This template set is generally considered to be a good compromise for AGB spectra since it captures the temporally extreme pulsations that in practice vary even more rapidly than stellar evolution isochrones can describe. The Lancon AGB spectra was extended beyond the original 5,100 – 24,900 Å range to cover the total range needed for ProGeny using a blackbody extension (at the appropriate temperature) at either end.
@@
In terms of parameter coverage, the AGB stellar spectra have the sparsest and least regular coverage, with only 14 templates covering the approximate temperature range 2,000 – 4,000 K. They also do not have a specified (i.e. known) surface gravity, so for the purposes of isochrone parameter interpolation they are placed at a fiducial value of -1 (as can be seen in Figure 3). The consequence of this is mitigated by the ‘base’ and ‘extend’ stellar spectra extending down to a surface gravity of 0, which covers the majority of AGB evolutionary phases in practice. It is only the most extreme TP-AGB pulsations that require the use of the empirical Lancon libraries provided here.
3.2.3 Hot Stars and White Dwarfs
@@
The final stellar spectrum variant that must be included for suitable isochrone coverage is ‘hot’ stars including stellar remnants (i.e. white dwarfs and neutron stars, although the latter are too dim to be relevant in a SSP). This includes most stars hotter than 50,000 K and ones with a surface gravity 6 $\text{cm}/\text{s}^{2}$ or greater, i.e. extremely compact and hot objects. It is notable from 3 that only the MIST isochrones have significant coverage inside this regime, i.e. the PARSEC and BaSTI isochrones do not use or require ‘hot’ stellar spectra. The hot remnant stellar spectra are important for producing at least some of the UV upturn for old stellar populations with ProGeny (see Figure 2), but it should be noted that there is still a great deal of uncertainty about what contributes to UV upturn features in elliptical galaxies. For instance, it is not well understood how dominant exposed core helium burning could be for producing UV photons, and the thorough modelling of binary evolution certainly has a significant effect (Stanway & Eldridge, 2018).
@@
The fiducial ‘hot’ template spectra comes from Werner et al. (2003) (and earlier works). This library is generated by a one-dimensional line-blanketed non-Local Thermodynamic Equilibrium (i.e. non-LTE) code, and is considered appropriate for use with white dwarf spectra in particular. It is a combination of the TMAP2 and TMAP3 Werner data available online at the Spanish Virtual Observatory101010svo2.cab.inta-csic.es/theory/newov2/index.php. The library used reaches up to 190,000 K, and covers the range of temperature and surface gravity for the majority of hot stars and white dwarf remnants observed on the MIST isochrones (temperature 20,000 – 190,000 K and logarithmic surface gravity 5 – 9. For this reason it works reasonably well for all hot phases (including helium burning) and remnants (including white dwarfs). The very coolest white dwarfs are not quite covered (below 20,000 K), but these barely exist within the current age of the Universe, and are such low luminosity that they contribute very little flux to any reasonable SSP.
We note that the PHOENIX Allard ‘extend’ template spectra are often used for representing reasonably hot (sub 70,000 K) stellar photospheres. Beyond 4,000 K the PHOENIX code used has some limited non-LTE approximations to better represent ‘hot’ spectra. In detail it is not identical to full non-LTE spectra, where we compared matching (in temperate and surface gravity) PHOENIX Allard and PoWR OB stars (as described in Hainich et al., 2019). Between 1,000 – 10,000 Å(where the differences are most significant) fluxes are within a few percent on average, with the biggest deviations for certain spectral features being tens of per cent. In practice these spectra are mostly used for very young (sub 10 Myr) stellar populations. For most integrated SED fitting work the impact of using PoWR OB stars for hotter spectra (rather than Allard) will be negligible, but for extreme bursts there could be biases in derived properties at the tens of per cent level. For the rest of this work (and the associated ProGeny II companion paper) we use the Allard template spectra without the use of an additional OB template library. In the future development of ProGeny we will investigate adding dedicated OB stellar spectra as a further option. But as with Wolf-Rayet stars, keen users can add extra OB star spectral templates as desired when running ProGeny locally.
@@ 3.3 Initial Mass Function
Figure 5: Comparison of the three different classic IMFs available as standard within ProGeny. In all cases 1 solar mass of stars have been formed by the IMF. The bottom panel shows the ration of number counts for each IMF relative to the Chabrier IMF. The Chabrier and Kroupa IMFs are in general very similar, as can be seen in the bottom panel showing the number counts ratio of Kroupa and Salpeter relative to Chabrier (red and blue lines respectively). Ch = Chabrier, Kr = Kroupa, Sp = Salpeter.
@@
In total, the pre-built ProGeny IMFs come in seven different flavours of IMF: Salpeter, Kroupa, Chabrier, age-evolving Kroupa, age-evolving Lacey, metallicity-evolving Kroupa, and metallically evolving Lacey. By following the included examples it is relatively easy for users to encode their own Universal IMFs, or even other evolving (either age or metallicity) variants. An obvious extension might be to parameterise the evolution via a property other than stellar population age. It often makes more sense to think of IMF evolution as being relative to cosmic age rather than age of a particular stellar population (which can be achieved by specifying the look-back time of observation in ProGeny), or perhaps relative to the metallicity enrichment history of the star formation event in question. The basic age-evolving IMFs available with ProGeny are shown in Figure 6, but note the exact implementation presented here is somewhat arbitrary — there is no well informed model for exactly how the IMF might evolve with age (or metallicity), so the existence of these functions is really to aid user experimentation (i.e. the default values should not be treated as especially well justified). Longer term work (beyond the scope of this paper), now that ProGeny exists with this degree of flexibility, will be to better justify if (and how) these evolving IMF forms can be used for fitting SEDs with ProSpect.
Figure 6: Comparison of the different age-evolving IMFs available as standard within ProGeny. In all cases 1 solar mass of stars have been formed by the IMF. The bottom panel shows the ration of number counts for each IMF relative to the Chabrier IMF. Ch = Chabrier, Kr = Kroupa, Lc = Lacey.
@@ 4 Comparisons and Discussion
Below we make some instructive comparisons of ProGeny. First internally (where we vary the isochrones, stellar spectra and IMF), and secondly to well know literature SSPs (that are mixture of similar and dissimilar SPLs).
4.1 Internal Low-Level Comparisons
@@
It is important to understand the role that the key components in ProGeny have when generating spectra at specific ages and metallicities (the core low-level output of any SSP). The most significant variations in spectral output come from the choice of isochrone (MIST, PARSEC, or BaSTI in the case of ProGeny), the stellar spectra used (be they empirical or theoretical) and the type of IMF (both fixed and evolving, in our case).
4.1.1 Impact of Isochrones
In Figure 7 we make a direct comparison of the three main isochrones initially available with ProGeny (MIST, PARSEC and BaSTI), where the other key characteristics of the SSPs are kept fixed (as per the Figure legend). The bottom sub-panels show the median absolute deviation (MAD) of the spectra in dex, where the red-shaded region indicates where the MAD is less than 0.2 dex (a classic threshold of SSP inferred stellar mass accuracy, see Taylor et al., 2011; Robotham et al., 2020). We note that here we are using the isochrones as provided, without additional attempt at remedy or treatment for ‘missing’ physics.
Figure 7: Comparison of Z spectra at different stellar population ages for different isochrones. The ages presented are log spaced with 1 dex intervals, covering 10 Myr through to 10 Gyr to allow comparison between all isochrones initially available with ProGeny. The vertical dashed lines indicate the $u$ and Ks band central wavelengths, common limits for SED analysis work.
In general, except for the very youngest stellar populations, the spectra agree much better than 0.2 dex through the optical and NIR regime. The biggest tension in the NIR is clearly for the youngest (10 Myr) sample, where MIST and PARSEC produces enhanced NIR flux compared to BaSTI. Investigating the isochrones, it is clear this is a feature of MIST/PARSEC producing significant pre main-sequence stellar light, which is dominated by cool collapsing photospheres that peak in the NIR. Beyond this feature the generated SSPs tend to be most consistent at $\sim$1 micron.
In the UV the agreement is much poorer for all ages, with very different implied spectra. As discussed earlier in Section 3.1, this is entirely due to the physics incorporated within the isochrone regarding a mixture of remnant stars and horizontal branch stars. It is worth emphasising that the serious disagreement in predicted UV flux occurs short-wards of FUV filter ( 1,500 Å), and for typical SED analysis at low redshift this should not be a concern. When analysing higher redshift sources some care should be taken when including data in this rest-frame regime: it is likely including such observations to an SED fit does more harm than good given the lack of predictive agreement seen here. Note we are not even exploring the additional impact of binary evolution on the UV spectra, which can also have a dramatic impact (Eldridge et al., 2017).
4.1.2 Impact of Stellar Spectra
Similar to the above, in Figure 8 we make a direct comparison of the four main ‘base’ stellar spectra initially available with ProGeny (PHOENIX (Husser), C3K, MILES and BaSeL), where the other key characteristics of the SSPs are kept fixed (as per the Figure legend).
Figure 8: Comparison of Z metallicity spectra at different stellar population ages for different stellar spectral libraries. The ages presented are log spaced with 1 dex intervals, covering 10 Myr through to 10 Gyr to allow comparison between all stellar spectra initially available with ProGeny. The vertical dashed lines indicate the $u$ and Ks band central wavelengths, common limits for SED analysis work.
Agreement here is much better than that noted between isochrones, with much less the 0.2 dex MAD between the spectral predictions. The only slight discrepancy of note is for youngest extreme UV population (below 1,000 Å). The main advantages and disadvantages between the different options for stellar spectra are really native wavelength coverage, spectral resolution and atmospheric coverage (i.e. the details presented in Table 3). For broadband SED analysis, especially in the optical – NIR regime, the choice of stellar spectra has almost no tangible impact within ProGeny. Because of its dense atmospheric parameter sampling and good temperature coverage, we suggest a weak preference for the C3K stellar spectral library for use within ProGeny, but we note the PHOENIX (Husser) library has the potential for much high spectral resolution SSPs (0.01 Å sampling versus 1 Å in C3K).
4.1.3 Impact of Initial Mass Function
Finally, in Figure 9 we make a direct comparison of the three classic IMFs available with ProGeny (Kroupa, Chabrier and Salpeter), where the other key characteristics of the SSPs are kept fixed (as per the Figure legend).
Figure 9: Comparison of Z spectra at different stellar population ages for different IMFs. The ages presented are log spaced with 1 dex intervals, covering 10 Myr through to 10 Gyr to allow comparison between all stellar spectra initially available with ProGeny. The vertical dashed lines indicate the $u$ and Ks band central wavelengths, common limits for SED analysis work.
The differences seen here are negligible, bar the general reduction in spectral normalisation seen for Salpeter. This is a direct consequence of the Salpeter IMF putting proportionally more mass into low luminosity stars. These stars are sufficiently sub-luminous that they have little impact on the spectral shape, so the overall effect is a drop in flux density (i.e. an increase in mass-to-light ratio).
4.1.4 Summary of Low-Level SSP Impact
From the above analysis it is clear that within a SPL, holding the methodology of how ingredients are combined fixed, the relative impact of isochrones, stellar spectra and IMFs have a clear hierarchy (within the bound of the IMF being a ‘classic’ local Universe variety, not an exotic type). The choice of isochrone has by far the most impact throughout the spectrum, with particularly notable uncertainty in the extreme UV. In this regime various aspects of complicated physics (including, but not limited to, remnant stars, horizontal branch stars and planetary nebulae) become dominant. In some SPLs interventions are made to incorporate some of these physical effects, but ProGeny relies on the available physics of the isochrones only. Beyond the choice of isochrones, the choice of stellar spectra has the next most impact. The main driver of selecting which one we prefer in ProGeny is the atmospheric parameter coverage, wavelength coverage and resolution — in general for similar atmospheric parameters the stellar spectra are highly consistent, with only small deviations in the UV and IR. Finally, the choice of IMF has little impact on the spectral shape, but some notable impact on the normalisation (mass-to-light), where the impact of this is discussed in more detail in the next Section. These general findings are consistent with the literature consensus discussed in the Introduction (Charlot et al., 1996; Maraston, 2005; Martins & Coelho, 2007; Percival & Salaris, 2009; Conroy et al., 2009), but ProGeny offers an easy route for interested researchers to explore such details (and more subtle behaviour) via its open software and detailed scripts and vignettes.
4.2 Internal High-Level Comparisons